autonomous consumption c1=R200
autonomous investment =R600
marginal propensity to consume c= 3/4
the equilibrium level of income is
(1) R8000
(2) R32000
(3) R6000
(4) R2000
1
Expert's answer
2019-09-19T09:29:44-0400

Y=C+I

C=a+cY

Y=200+0.75Y+600

0.25Y=800

Y=3200

Since this answer doesn't exist, I've decided that it is 8000